OpenStudy (smurfy14):

Is my answer correct? Find the equation of the line in slope intercept form for a line through points (-1,-4) and (3,2). I got y=1/2(3x-5)

OpenStudy (anonymous):

WHICH CLRETRICE ?

OpenStudy (kira_yamato):

If we want it in the "y=mx+c" form, then find the gradient and use the formula y-y0 = m(x - x0). Your answer will be y = mx + (y0 - x0)

OpenStudy (kira_yamato):

Use any one of the 2 pts given for (x0, y0)